a stone is dropped from a top of building covers 36% of total distance in its last second of its motion. the height of building is ?

Grade:11

1 Answers

Arun
6 years ago

We can write 36/100 = 9/25 also

Let ‘n’ be the time of fall.

Using, S = ut + ½ at2

=> h = 0 + ½ gn2

=> n = [(2h)/g]1/2

Distance travelled in the nth second is,

Sn = u + (a/2)(2n – 1)

=> 9h/25 = 0 + (9.8/2)[2{(2h)/g}1/2 – 1]

=> 9h/25 = 4.9[2{2h/9.8}1/2 – 1]

Solving for h we get,

h = 122.5 m